A waterbed heater uses 450 W of power. It is on 35% of the time, off 65%.
What is the annual cost of electricity at a billing rate of $0.11kW/h ?
Answer #1
The power of the heater P = 450W
The time of heater on 35%
The rate of billing $0.11kWh
For annual cost
1year = 365*24hr = 8760hr
But it is on only 35% so
time = 8760*0.35 = 3066hr
The Energy consumed
Q = P*t = 450W (3066hr)
= 1379700 J
= 1379.7kWh
The electricity bill
= 1379.7kWh(0.11kWh) = $151.76